# Conjugacy properties of multivariate unified skew–elliptical distributions

### v2

@@ 4 Conclusions
This article proves that sue distributions possess important conjugacy properties when combined with broad classes of likelihoods that generalize classical probit, tobit, multinomial probit, and linear models in several directions. These generalizations include multivariate representations based on general elliptical noise terms and allow for asymmetric representation relying on unified skew–elliptical extensions. Our conjugacy results leverage available and newly–derived closure properties of the sue family to prove that priors within such a class yield again sue posterior distributions when combined with the likelihood of the models mentioned above, under the classical Bayes rule. Recalling Propositions 13, these results are technically derived by starting from a joint sue distribution for the parameters and the observed data. Such a proof technique is not meant to provide a different perspective on the standard specification of a prior and a likelihood in Bayesian statistics. Rather, it provides a convenient strategy that facilitates the derivation, within the sue class, of meaningful priors and likelihoods yielding closed–form sue posterior distributions.
@@
More specifically, Examples 16 clarify that our results include Bayesian models of direct interest in practice, such as those based on multivariate Gaussian or Student’s $t$ formulations, along with the corresponding skewed extensions. In this respect, an interesting direction would be to specialize Propositions 13 to other sue sub–families, such as those based on Cauchy or logistic density generators. This can be accomplished by replacing the generic density generators in Propositions 13, with those yielding the sub–family investigated. These extensions further motivate advancements in the study of other relevant sue sub–families to derive results and properties similar to those characterizing sun [e.g., 6] and sut [e.g., 62] distributions. Particularly impactful, within our context, would be the derivation of additive stochastic representations as those obtained for suns and suts. Advancements along these lines would facilitate i.i.d. sampling under any sue posterior distribution, thereby enlarging the class of models and priors that allow for tractable Bayesian inference. The recent additive stochastic representations derived by Yin and Balakrishnan [63] for general skew–elliptical distributions provide a promising advancement in this direction, which also suggests that related results could be derived even for the wider sue family. Similarly, expanding the available strategies for the efficient evaluation of the moments of sue distributions in, e.g., (6)–(7), would further facilitate Bayesian inference leveraging the conjugacy results derived in the present article. The contributions by [e.g., 13, 54, 39, 53, 59] provide important results along these lines which motivate future research to showcase the computational advantages and the practical impact of these solutions when the focus is on Bayesian inference under the newly–derived sue posterior distributions.
Finally, we shall emphasize that the sue family can be itself rephrased as a particular case of selection elliptical distributions [7] arising from even more general conditioning mechanisms. As such, it would be interesting to expand the conjugacy properties derived in this article for sue distributions to the broader selection elliptical family. This is expected to further enlarge the class of models and priors admitting closed–form posterior distributions.
@@
We shall conclude by highlighting that not all the priors and likelihoods implied by the general results in Propositions 13 have direct practical applicability. Nonetheless, from a theoretical perspective, also these instances are of interest in expanding the analysis of the probabilistic properties of the sue family. Moreover, although conjugacy is a desirable property, it is important to emphasize that those priors and likelihoods in the sue family that do not yield sue posteriors, can still allow for Bayesian inference leveraging, e.g., mcmc methods or deterministic approximations.

### v3

@@ 4 Conclusions
This article proves that SUE distributions have important conjugacy properties when combined with broad classes of likelihoods that generalize classical probit, tobit, multinomial probit, and linear models in several directions. These generalizations include multivariate representations based on general elliptical noise terms and allow for asymmetric representation relying on unified skew-elliptical extensions. Our results leverage available and newly-derived closure properties of the SUE family to prove that priors within this class yield again SUE posterior distributions when combined with the likelihood of the models mentioned above, under the classical Bayes rule. Recalling Propositions 13, these results are technically derived by starting from a joint SUE distribution for the parameters and the observed data. Such a proof technique is not meant to provide a different perspective on the standard specification of a prior and a likelihood in Bayesian statistics. Rather, it provides a convenient strategy that facilitates the derivation, within the SUE class, of meaningful priors and likelihoods yielding closed-form SUE posterior distributions.
@@
More specifically, Examples 16 clarify that our results include Bayesian models of direct interest in practice, such as those based on multivariate Gaussian or Student’s $t$ formulations, along with the corresponding skewed extensions. In this respect, an interesting direction would be to specialize Propositions 13 to other SUE sub-families, e.g., those based on Cauchy or logistic density generators. This can be accomplished by replacing the generic density generators in Propositions 13, with those yielding the sub-family investigated. These extensions further motivate advancements in the study of other relevant SUE sub-families to derive results and properties similar to those characterizing SUN [e.g., 6] and SUT [e.g., 62] distributions. Particularly impactful, within our context, would be the derivation of additive stochastic representations as those obtained for SUNs and SUTs. Advancements along these lines would facilitate i.i.d. sampling under any SUE posterior, thus enlarging the class of models and priors that allow for tractable Bayesian inference. The recent additive stochastic representations derived by Yin and Balakrishnan [63] for general skew-elliptical distributions provide a promising advancement in this direction, which also suggests that related results could be derived even for the wider SUE family. Similarly, expanding the available strategies for the efficient evaluation of the moments of SUE distributions in, e.g., (6)–(7), would further facilitate Bayesian inference leveraging the conjugacy results derived in the present article. Current contributions [e.g., 13, 54, 39, 53, 59] provide important results along these lines which motivate future research to showcase the computational advantages and the practical impact of these solutions when the focus is on Bayesian inference under the newly-derived SUE posterior distributions.
@@
Finally, we shall emphasize that the SUE family can be itself rephrased as a particular case of selection elliptical distributions [7] arising from even more general conditioning mechanisms. As such, it would be interesting to expand the conjugacy properties derived in this article for SUE distributions to the broader selection elliptical family. Let us conclude by highlighting that not all the priors and likelihoods implied by the general results in Propositions 13 have direct practical applicability. Nonetheless, from a theoretical perspective, also these instances are of interest in expanding the analysis of the probabilistic properties of the SUE family. Moreover, although conjugacy is a desirable property, it is important to emphasize that those priors and likelihoods in the SUE family that do not yield SUE posteriors, can still allow for Bayesian inference leveraging, e.g., MCMC methods and deterministic approximations.
Acknowledgments
Maicon J. Karling and Marc G. Genton acknowledge the support by the King Abdullah University of Science and Technology (KAUST) in Saudi Arabia.